OpenStudy (anonymous):

Find the slope and the y-intercept of the equation y − 3(x − 1) = 0

OpenStudy (xapproachesinfinity):

to find the slope intercept form solve for y \(\huge \rm y-3(x-1)=0\Longrightarrow y-3x+3=0\) \(\huge \rm y-3x=-3\Longrightarrow y=3x-3\)

OpenStudy (xapproachesinfinity):

\(\huge \rm recall~y-intercept~form ~is~ y=mx+b\) where m is the slope and (0,b) is the y intercept

OpenStudy (xapproachesinfinity):

the part that is not appearing is y=mx+b

OpenStudy (anonymous):

ok it ok thank you

OpenStudy (xapproachesinfinity):

good luck! try playing with this so you know what you are doing!